The numbers of absent days a student had in the last 5 months are {0,4,2,0,3}.
What is the approximate variance of this data?

my options are:
F 1.6
H 2.6
G 1.8
J 2.8

Answer by edjones(8007)

m=9/5=1.8
.
0-1.8=-1.8^2=3.24
4-1.8=2.2^2=4.84
2-1.8=0.2^2=.04
0-1.8=-1.8^2=3.24
3-1.8=1.2^2=1.44
===========>12.8
.
12.8/4=3.2 Variance
.
Ed